OpenStudy (anonymous):

Explain, in complete sentences, the effect of the difference of squares pattern on the multiplication of radicals. Give an example.

terenzreignz (terenzreignz):

If, for instance, we have \[a + \sqrt{b}\] and we multiply it with its conjugate \[a - \sqrt{b}\] where a is a real number and b is a positive number which is not a perfect square, then all we need remember is the pattern of "difference of two squares", like so : \[(x + y)(x - y) = x ^{2} - y^{2}\] In our case, \[x = a\] and \[y = \sqrt{b}\] So the result of multiplying them is \[a ^{2} - (\sqrt{b})^{2} = a^{2} - b\] In other words, the effect of the "difference of two squares" pattern is the loss of the radical sign, typically used to simplify equations. For example \[(2 + \sqrt{3})(2 - \sqrt{3}) = 2^{2} - (\sqrt{3})^{2} = 4 - 3 = 1\]